Find the law of motion of a particle in the field U(x) = – Bx2, if its total energy is zero (B is positive constant). 

1
Expert's answer
2021-06-15T16:16:59-0400
"F=\frac{dU}{dx}=-2Bx\\ma=-2Bx\\\omega=\sqrt{\frac{2B}{m}}\\x=A\cos{(\omega t+\phi)}"